OpenStudy (anonymous):

Assume that the number of defective basketballs produced is related by a linear equation to the total number produced. Suppose that 6 defective balls are produced in a lot of 325, and12 defective balls are produced in a lot of 525. Find the number of defective balls produced in a lot of 775 balls.____________?

OpenStudy (anonymous):

we have unlimited attempts

Directrix (directrix):

My take on this is to write an equation joining (6,325) and (12, 525). That equation will predict the number y of balls in a given lot where x is the number of defective balls. y = (100/3) x + 125 775 = (100/3) x + 125 x = 39/2 = 19.5 balls which is 20 (approx) I hope somebody comes by and checks this.

OpenStudy (anonymous):

its correct

OpenStudy (anonymous):

I have posted some others thank you very much